OpenStudy (anonymous):

how many years will it take for a $50 investment to grow to $5,000 if it is compounded continuously at a rate of 5%? Round your answer to 2 decimal places.

OpenStudy (anonymous):

I got 92.10 ???

OpenStudy (blurbendy):

5000 = 50(e^(.05t)) 500/50 = e^(.05t) ln(100) ------ = t .05

OpenStudy (anonymous):

so my answer is correct?

OpenStudy (blurbendy):

yup, good job

OpenStudy (anonymous):

k thanks :)

OpenStudy (blurbendy):

yw